Given that f(x) = 4x + 3 and g(x)=x*, find (fog)(-4). (fog)(-4)=
Answer #1

solution f(x) = 4x+3 ; gox)=x gra) (fog) (-4) = f (g (-4)) no tu put a=-4 in af (C-43 ) =f(-64) van put x=-64 = 46-64) +3 in